### v1

# OneForecast: A Universal Framework for Global and Regional Weather Forecasting

Abstract
Accurate weather forecasts are important for disaster prevention, agricultural planning, and water resource management. Traditional numerical weather prediction (NWP) methods offer physically interpretable high-accuracy predictions but are computationally expensive and fail to fully leverage rapidly growing historical data. In recent years, deep learning methods have made significant progress in weather forecasting, but challenges remain, such as balancing global and regional high-resolution forecasts, excessive smoothing in extreme event predictions, and insufficient dynamic system modeling. To address these issues, this paper proposes a global-regional nested weather forecasting framework based on graph neural networks (GNNs). By combining a dynamic system perspective with multi-grid theory, we construct a multi-scale graph structure and densify the target region to capture local high-frequency features. We introduce an adaptive information propagation mechanism, using dynamic gating units to deeply integrate node and edge features for more accurate extreme event forecasting. For high-resolution regional forecasts, we propose a neural nested grid method to mitigate boundary information loss. Experimental results show that the proposed method performs excellently across global to regional scales and short-term to long-term forecasts, especially in extreme event predictions (e.g., typhoons), significantly improving forecast accuracy. Our codes are available at https://github.com/YuanGao-YG/OneForecast.
Machine Learning, ICML
1 Introduction
Accurate weather forecasting is crucial for disaster prevention, optimizing agriculture and energy planning, and ensuring water resource management (Chen et al., 2023a; Pathak et al., 2022; Ahmed et al., 2021). Traditional Numerical Weather Prediction (NWP) methods (Bauer et al., 2015) rely on the numerical solution of atmospheric dynamic equations (Achatz et al., 2023; Buzzicotti et al., 2023), ensuring consistency across spatiotemporal scales from a physical perspective. However, with the growing volume of observational and historical data, and the increasing demand for high-resolution and long-term forecasts, NWP methods often struggle with computational costs and fail to fully leverage the potential value of vast data.
Figure 1: Forecast results of extreme cyclones. (a) OneForecast’s predicted wind speed for Typhoon Molva (2020) at 850 hPa pressure level with a 60-hour lead time. (b)–(c) the predicted cyclone tracks of Typhoon Yagi (2018) and Typhoon Molva (2020) using different models
Recent developments in deep learning (DL) methods offer new perspectives for weather forecasting. Early spatio-temporal prediction algorithms (Wu et al., 2024a), such as ConvLSTM (Shi et al., 2015) and PredRNN (Wang et al., 2022), focus on regional precipitation. Recent large-scale scientific computing models, like Pangu-weather (Bi et al., 2023), GraphCast (Lam et al., 2023), and NowcastNet (Zhang et al., 2023), achieve significant results in medium- and short-term forecasts and show high potential in extreme event prediction (e.g., precipitation and Typhoon track prediction) (Chen et al., 2024; Espeholt et al., 2022; Gong et al., 2024; Wu et al., 2024b). However, pure AI methods still face several core challenges:
❶ Global and regional high-resolution forecasts are hard to balance. Regional predictions often lack boundary information, making it difficult to effectively nest global data. ❷ Extreme events and long-term forecasts suffer from over-smoothing. They fail to capture high-frequency disturbances, leading to reduced forecast accuracy. ❸ Lack of dynamic system modeling capability. This is especially true for capturing complex interactions between nodes at multiple scales and learning high-frequency node-edge features.
To address these challenges, we propose OneForecast, a global-regional nested weather forecasting framework based on Graph Neural Networks (GNNs). Inspired by heuristic learning from numerical methods, we construct a multi-scale graph structure based on dynamical systems and multi-grid theory (He & Xu, 2019; He et al., ), refining it for the target region to capture local high-frequency features with greater detail. Additionally, to solve issues like over-smoothing in extreme events and long-term forecasts, which hinder the capture of high-frequency disturbances, we introduce an adaptive information propagation mechanism. This mechanism deepens the integration of node and edge features through dynamic gating units. Finally, for regional high-resolution forecasting, we adopt a nested grid strategy (Phillips & Shukla, 1973) that inherits large-scale background information from the global scale, significantly alleviating the boundary information loss. Through this integrated framework, we aim to effectively capture high-frequency features and extreme events across global to regional scales, as well as from short-term to long-term forecasts.
The method most similar to ours is Graph-EFM (Oskarsson et al., 2024). It also uses a hierarchical graph neural network for global and regional weather modeling. However, for high-resolution regional forecasts, it treats global low-resolution data as non-trainable forcing conditions, making it unable to adaptively couple multi-scale information based on actual needs. And it doesn’t treats the forecasts of the global model in the region as forcing, which unable to fully utilize the information of the global model. In contrast, our neural nested grid method applies trainable local refinement to the target region in the network structure. It also updates boundary and background information with global model future forcing dynamically through end-to-end training during global-regional coupling. This design better captures the interaction between large-scale global backgrounds and high-frequency regional details. Our experiments (Sec 3.3) show that our method achieves greater stability and accuracy in long-term rolling inference.
The contribution of this paper can be summarized as follows: (1) Global-Regional Unified Forecasting Framework. We propose a Graph Neural Network method that supports both global scale and regional high-resolution forecasting, achieving high-accuracy results for multi-scale and multi-time frame weather forecasts within the same framework. (2) Adaptive Information Propagation Mechanism. Through Dynamic Gating Units and graph attention modules, we deeply integrate node and edge features, more accurately capturing extreme events and other high-frequency disturbance signals within the multi-scale graph structure. As shown in Figure 1, OneForecast delivers better performance in tracking extreme events like typhoons. (3) Nested Grid and Long-Term Forecasting. By using a nested grid to merge global and regional information, we overcome the boundary loss issue in regional forecasting. This method effectively mitigates the loss of details caused by over-smoothing in long-term forecasts.
2 Method
Figure 2: Overview of Our OneForecast. (a) The overall architecture includes input variables, an encoder, a message passing module, a decoder, and visualization of forecast variables; (b) The global forecasts module uses rollout technology to generate future forecasts; (c) The neural nested grid method specializes in regional high-resolution weather forecasts tasks; and (d) The ensemble forecasting module generates long-term forecast results.
Problem Definition In this study, we model weather forecasting as an autoregressive problem (Lam et al., 2023). At each time step $t$, we use the meteorological state comprising surface variables $\mathbf{X}_{t}$ and pressure level variables $\mathbf{P}_{t}$ to forecast the state at the next time step. We concatenate the surface and pressure level variables along the channel dimension to form the combined input: $\mathbf{Z}_{t}=[\mathbf{X}_{t},\mathbf{P}_{t}]\in\mathbb{R}^{N\times d}$, where $N=H\times W$ represents the number of grid locations (nodes), and $d=d_{x}+d_{p}$ is the total number of variables. Here, $d_{x}$ and $d_{p}$ are the numbers of surface level and pressure level variables, respectively. In our setup, the initial input contains 69 variables: 4 surface level variables and 65 pressure level variables. Our model aims to forecast the combined variables at the next time step $\hat{\mathbf{Z}}_{t+1}$ using the current input $\mathbf{Z}_{t}$, capturing the spatiotemporal evolution of the atmosphere: $\hat{\mathbf{Z}}_{t+1}=\text{Model}(\mathbf{Z}_{t};\Theta),$ where $\Theta$ denotes the model parameters. The training objective is to minimize the relative mean squared error (MSE) between the forecasts and the true values across all time steps: $\min_{\Theta}\frac{1}{T}\sum_{t=0}^{T-1}\frac{\left\|\hat{\mathbf{Z}}_{t+1}-\mathbf{Z}_{t+1}\right\|_{2}^{2}}{\left\|\mathbf{Z}_{t+1}\right\|_{2}^{2}}$. During inference, we adopt a rollout strategy to forecast longer sequences. Starting from the initial state $\mathbf{Z}_{0}$, the model recursively uses its previous forecasts as the next input: $\hat{\mathbf{Z}}_{t+1}=\text{Model}(\hat{\mathbf{Z}}_{t};\Theta),\quad t=0,1,2,\dots,T-1.$ This strategy allows the model to generate extended weather forecasts using its own forecasts.
2.1 Earth-specific Region Refined Graph Encoder
In the encoder of OneForecast, inspired by (Fortunato et al., 2022)  (Lam et al., 2023), we introduce an Earth-specific Region Refined Multi-scale Graph to improve the interaction of node features in complex dynamical systems. Inspired by the idea of multigrid methods (He & Xu, 2019), we construct a multi-level Graph Neural Network architecture that includes grids of multiple granularities. Each grid has the same number of nodes but different grid densities, thereby capturing spatial features at different scales. Specifically, we define the multi-scale graph structure as:
$\mathcal{G}=\left(\mathcal{V}^{G},\mathcal{V},\mathcal{E}^{(1)},\mathcal{E}^{(2)},\ldots,\mathcal{E}^{(L)},\mathcal{E}^{(R)},\mathcal{E}^{\mathrm{G}2\mathrm{M}},\mathcal{E}^{\mathrm{M}2\mathrm{G}}\right),$ (1)
where ${\mathcal{V}}^{G}$ represents the set of lat-lon grid nodes, with a total of $N=H\times W$ nodes; $\mathcal{V}$ represents the mesh nodes. $\mathcal{E}^{(l)}$ denotes the edge set at the $l$-th scale, corresponding to grids of different granularities, where $l=1,2,\dots,L$, and $\mathcal{E}^{(R)}$ represents regional refined edges. $\mathcal{E}^{\mathrm{G}2\mathrm{M}}$ and $\mathcal{E}^{\mathrm{M}2\mathrm{G}}$ are the unidirectional edges that connect lat-lon grid nodes and mesh nodes. All scales share the same set of nodes $\mathcal{V}$. More details can be found in Appendix D.
In the encoder, we first map the input meteorological state $\mathbf{Z}_{t}\in\mathbb{R}^{N\times d}$ to the initial node feature representation:
$\mathbf{h}_{i}^{(0)}=\phi(\mathbf{Z}_{t,i}),\quad i=1,2,\ldots,N,$ (2)
where $\phi(\cdot)$ is the feature mapping function, and $\mathbf{Z}_{t,i}$ is the input feature at node $i$. Next, we iteratively update the node features on the multi-scale graph structure. At iteration $k$, the feature update formula for node $i$ is:
$\mathbf{h}_{i}^{(k)}=\sigma\left(\sum_{l=1}^{L}\sum_{j\in\mathcal{N}_{i}^{(l)}}\mathbf{W}^{(l)}\mathbf{h}_{j}^{(k-1)}+\mathbf{b}^{(l)}\right),$ (3)
where $\mathcal{N}_{i}^{(l)}$ is the set of nodes adjacent to node $i$ at the $l$-th scale, $\mathbf{W}^{(l)}$ is the weight matrix at the $l$-th scale, $\mathbf{b}$ is the bias term, and $\sigma(\cdot)$ is the activation function. To enhance the forecasting accuracy in specific regions, we introduce a region-refined grid on the finest global grid. For nodes within the target region, we add denser edge connections to capture local high-frequency features. In this way, the update of node features not only considers global multi-scale information but also incorporates region-specific fine-grained information.
2.2 Multi-stream Messaging
To address the issue of information transmission between nodes in complex dynamic systems, we propose a module called Multi-stream Messaging (MSM). This module consists of an adaptive messaging mechanism, including a dynamic multi-head gated edge update module and a multi-head node attention mechanism. And OneForecast includes 16 MSMs for messaging.
Dynamic Multi-head Gated Edge Update Module. Unlike traditional message passing methods based on MLPs, we introduce dynamic gating and multi-head mechanisms to control the information flow more precisely. For each edge, we concatenate its own features with those of the source node and the target node:
$\mathbf{c}_{i}=\operatorname{Concat}\left(\mathbf{e}_{i},\mathbf{h}_{s(i)},\mathbf{h}_{d(i)}\right)\in\mathbb{R}^{D_{e}+2D_{h}},$ (4)
where $\mathbf{e}_{i}$ is the feature of edge $i$, $\mathbf{h}_{s(i)}$ and $\mathbf{h}_{d(i)}$ are the features of the source and target nodes of edge $i$, $D_{e}$ is the edge feature dimension, and $D_{h}$ is the node feature dimension. Next, we generate gating vectors through a two-layer MLP to regulate the information flow. Specifically, the gating vector is divided into three parts: edge feature update gate $g_{e}$, source node feature gate $g_{s}$, and destination node feature gate $g_{d}$. First, we perform the first layer linear transformation and activation:
$\mathbf{z}_{i}=\operatorname{SiLU}\left(\mathbf{W}_{1}\mathbf{c}_{i}+\mathbf{b}_{1}\right),$ (5)
where $\mathbf{W}_{1}\in\mathbb{R}^{h\times(D_{e}+2D_{h})}$ is the weight matrix of the first layer, and $\mathbf{b}_{1}\in\mathbb{R}^{h}$ is the bias term. Then, we perform the second layer linear transformation and Sigmoid activation:
$\mathbf{g}_{i}=\sigma\left(\mathbf{W}_{2}\mathbf{z}_{i}+\mathbf{b}_{2}\right)\in\mathbb{R}^{3HD},$ (6)
where, $\mathbf{W}_{2}\in\mathbb{R}^{3HD\times h}$ is the weight matrix of the linear transformation of the second layer, and $D$ is the feature dimension for each gate. For each head $h$, the gating values $\mathbf{g}_{i}^{(h,e)}$, $\mathbf{g}_{i}^{(h,s)}$, and $\mathbf{g}_{i}^{(h,d)}$ are vectors of dimension $D$, corresponding to the edge feature gate, source node feature gate, and destination node feature gate, respectively.
Subsequently, we use Edge Sum MLP (ESMLP) to perform linear transformation and nonlinear activation on the edge features to generate the updated edge features:
$\mathbf{e}_{i}^{\prime}=\operatorname{ESMLP}_{e}\left(\mathbf{e}_{i},\mathbf{h}_{s(i)},\mathbf{h}_{d(i)}\right)\in\mathbb{R}^{D_{e}^{\prime}},$ (7)
where $D_{e}^{\prime}$ is the dimension of the updated edge features. Finally, we combine the gating vectors and the updated edge features to generate the final updated edge features through weighted averaging and residual connections:
$\mathbf{e}_{i}^{\text{new}}=$ $\frac{1}{3}\sum_{h=1}^{H}\bigg{(}\mathbf{g}_{i}^{(h,e)}\odot\mathbf{e}_{i}^{\prime}+\mathbf{g}_{i}^{(h,s)}\odot\mathbf{h}_{s(i)}$ (8)
$\quad+\mathbf{g}_{i}^{(h,d)}\odot\mathbf{h}_{d(i)}\bigg{)}+\mathbf{e}_{i},$
where $\odot$ denotes element-wise multiplication.
Multi-head Node Attention Mechanism. Compared to traditional message passing mechanisms, multi-head attention mechanisms can more precisely capture complex dependencies between nodes and dynamically adjust the way information is aggregated through attention weights. For each edge $e_{i}=(j\rightarrow k)$, we use an MLP to calculate the attention score:
$\mathbf{a}_{i}=\operatorname{MLP}_{a}\left(\mathbf{e}_{i}^{\text{new}}\right)\in\mathbb{R}^{H},$ (9)
then, we normalize the attention scores:
$\alpha_{i}^{(h)}=\frac{\exp\left(\mathbf{a}_{i}^{(h)}\right)}{\sum_{e_{j}\in\mathcal{E}(k)}\exp\left(\mathbf{a}_{j}^{(h)}\right)},\quad\forall h=1,2,\dots,H,$ (10)
where $\mathcal{E}(k)$ denotes the set of all incoming edges to node $k$, and $\alpha_{i}^{(h)}$ is the attention weight of the $h$-th attention head for edge $e_{i}$. Next, we perform weighted aggregation of the edge features. For each node $k$, based on the attention weights, we compute the weighted sum of the features of all edges incoming to node $k$, generating the aggregated feature for each head:
$\mathbf{m}_{k}^{(h)}=\sum_{e_{i}\in\mathcal{E}(k)}\alpha_{i}^{(h)}\cdot\mathbf{e}_{i}^{\text{new}}\in\mathbb{R}^{D_{e}^{\prime}},$ (11)
then, we flatten and concatenate the aggregated features from all heads:
$\mathbf{M}_{k}=\operatorname{Flatten}\left[\mathbf{m}_{k}^{(1)},\mathbf{m}_{k}^{(2)},\dots,\mathbf{m}_{k}^{(H)}\right]\in\mathbb{R}^{D_{e}^{\prime}\cdot H}.$ (12)
Finally, we concatenate the aggregated edge features with the original node features and, through an MLP, perform a nonlinear transformation to generate the updated node features:
$\mathbf{h}_{k}^{\text{new}}=\operatorname{MLP}_{n}\left(\operatorname{Concat}\left(\mathbf{M}_{k},\mathbf{h}_{k}\right)\right)+\mathbf{h}_{k}.$ (13)
In summary, in each iteration, we use the multi-stream messaging module to update the node features. Specifically, the node feature update formula is:
$\mathbf{h}_{i}^{(k)}=\sigma\left(\sum_{l=1}^{L}\operatorname{MSM}\left(\mathbf{h}_{i}^{(k-1)},\mathcal{E}^{(l)}\right)+\mathbf{b}\right),$ (14)
where $\operatorname{MSM}$ represents the aforementioned multi-stream messaging operation, $\mathcal{E}^{(l)}$ is the set of edges at the $l$-th scale, and $\sigma(\cdot)$ is the activation function. In the region-refined graph structure, for nodes within the target region, we additionally consider the set of edges within the region $\mathcal{E}^{\text{region}}$ to capture finer local information.
Theoretical Analysis. From a theoretical perspective, we explain why our method helps capture high-frequency information. This enhances long-term prediction ability and improves the ability to detect extreme events.
Theorem 2.1.
High-pass Filtering Property of Multi-stream Messaging. Considering the improved multi-stream message passing mechanism, suppose the graph signal $\bm{f}\in\mathbb{R}^{N}$ has a spectrum $\hat{\bm{f}}=\bm{U}^{\top}\bm{f}$ under the graph Fourier basis $\bm{U}=[\bm{u}_{1},...,\bm{u}_{N}]$, where $\bm{L}=\bm{U}\bm{\Lambda}\bm{U}^{\top}$ is the normalized graph Laplacian matrix and $\bm{\Lambda}=\operatorname{diag}(\lambda_{1},...,\lambda_{N})$ is its eigenvalue diagonal matrix ($0\leq\lambda_{1}\leq...\leq\lambda_{N}\leq 2$). Define the frequency response function of the message passing operator as $\rho:\lambda\mapsto\mathbb{R}$. If the dynamic gating weights satisfy:
$g^{(h,e)}_{i},g^{(h,s)}_{i},g^{(h,d)}_{i}\propto|\lambda_{i}-1|+\epsilon\quad(\epsilon>0)$ (15)
then there exist constants $\alpha>0$ and $\kappa>0$ such that the frequency response of the operator satisfies:
$\rho(\lambda_{i})\geq\alpha|\lambda_{i}-1|\quad\text{and}\quad\rho(\lambda_{i})\geq\kappa\lambda_{i}$ (16)
that is, the operator is a strictly high-pass filter.
The proof of Theorem A.1 can be found in Appendix A.
2.3 Decoding and Optimization
The decoder’s goal is to decode the latent information back to meteorological variables on the latitude-longitude grid. We obtain the updated feature representation $\mathbf{h}_{i}$ for each node. For each node $i$, the decoder applies the mapping function:
$\hat{\mathbf{Z}}_{t+1,i}=\psi(\mathbf{h}_{i}),$ (17)
where $\psi(\cdot)$ is an MLP that converts the latent node features into the predicted variables $\hat{\mathbf{Z}}_{t+1,i}$ for the next time step.
We use relative $L_{2}$ loss function for model training. The loss function is defined as:
$\mathcal{L}=\frac{1}{KHW}\sum_{k=1}^{K}\sum_{i=1}^{H}\sum_{j=1}^{W}\frac{\left(\hat{x}_{i,j,k}^{t+l\delta t}-x_{i,j,k}^{t+l\delta t}\right)^{2}}{\left(x_{i,j,k}^{t+l\delta t}\right)^{2}},$ (18)
where $\hat{x}_{i,j,k}^{t+l\delta t}$ and $x_{i,j,k}^{t+l\delta t}$ are the predicted and true values for variable (channel) $k$ at spatial location $(i,j)$ and time $t+l\delta t$; $K$ is the number of variables (channels); $H$ and $W$ are the height and width of the spatial dimensions, respectively; $\delta t$ is the time interval of single-step prediction (we use $\delta t=6$ hours).
2.4 Downstream Tasks
We consider three principal downstream tasks:
Global Weather Forecasting. As detailed in Section 2 and illustrated in Figure 2(b), we employ a rollout approach during inference, using the trained model for multi-step extrapolation. Specifically, starting from the initial state $\mathbf{Z}_{0}$, the model recursively uses its previous predictions as inputs for subsequent time steps, generating a sequence of future global weather forecasts.
Regional High-Resolution Forecasting. To enhance the accuracy of high-resolution forecasts in specific regions, we propose a neural nested grid method, illustrated in Figure 2(c). This method combines global low-resolution future forecasts with regional high-resolution data to produce detailed forecasts for the target region. We first input the global low-resolution data at time $t$ into the pre-trained global model to obtain the global forecasts $\hat{\mathbf{Z}}_{t+1}^{\text{global}}$ at time $t+1$. We extract $\hat{\mathbf{Z}}_{t+1}^{\text{global1}}$ from $\hat{\mathbf{Z}}_{t+1}^{\text{global}}$, which shares the same spatial range as the region, and $\hat{\mathbf{Z}}_{t+1}^{\text{global2}}$, which includes the boundary of the region (the boundary are defined as two grid points around the region). Both $\hat{\mathbf{Z}}_{t+1}^{\text{global1}}$ and $\hat{\mathbf{Z}}_{t+1}^{\text{global2}}$ are then interpolated to match the resolution of the high-resolution regional data, which are concatenate as $\hat{\mathbf{Z}}_{t+1}^{\text{global}}$ to acted as global force. We then combine the regional high-resolution data at time $t$ with the $\hat{\mathbf{Z}}_{t+1}^{\text{global}}$ to form the input to the regional model. The global forecasts provide the necessary boundary conditions for the regional forecasts. The regional model then produces the high-resolution forecasts for the regional state at time $t+1$:
$\small\hat{\mathbf{Z}}_{t+1}^{\text{region}}=\text{Model}_{\text{region}}\left(\operatorname{Concat}\left(\hat{\mathbf{Z}}_{t+1}^{\text{global}},\,\mathbf{Z}_{t}^{\text{region}}\right);\,\Theta_{\text{region}}\right).$ (19)
Long-Term and Esemble Weather Forecasting. The initial condition of the atmospheric state is uncertain, so reasonable quantification of this uncertainty is conducive to improve to forecast performance. In this work, we set N=50. To account for the uncertainty in the atmospheric initial state for long-term ensemble forecasting, we generate $N$ perturbed initial conditions $\mathbf{Z}_{0}^{(n)}$ by adding Perlin noise $\varepsilon^{(n)}$ to $\mathbf{Z}_{0}$ (Chen et al., 2023b). Each perturbed initial condition is input into the model, and through recursive rollout over $T$ time steps, we obtain individual forecasts $\hat{\mathbf{Z}}_{t+1}^{(n)}$. Finally, at each time step $t+1$, we compute the ensemble mean prediction $\hat{\mathbf{Z}}_{t+1}^{\text{ensemble}}=\frac{1}{N}\sum_{n=1}^{N}\hat{\mathbf{Z}}_{t+1}^{(n)}$ by averaging the forecasts from all $N$ ensemble members.
3 Experiments
Table 1: In all benchmark tests, we compare the performance of our study with 4 baselines. RMSE and ACC are recorded. A small RMSE ($\downarrow$) and a bigger ACC ($\uparrow$) indicate better performance. The best results are in bold, and the second best are with underline.
Model Metric
6-hour 1-day 4-day 7-day 10-day
RMSE ACC RMSE ACC RMSE ACC RMSE ACC RMSE ACC
Pangu-weather (Bi et al., 2023) 0.0921 0.9958 0.1635 0.9868 0.3723 0.9305 0.5576 0.8414 0.6680 0.7763
Fengwu (Chen et al., 2023a) 0.1428 0.9913 0.2584 0.9738 0.5186 0.8958 0.7874 0.8061 1.2370 0.7409
Graphcast (Lam et al., 2023) 0.0887 0.9962 0.1634 0.9878 0.3962 0.9308 0.6026 0.8362 0.7374 0.7625
Fuxi (Chen et al., 2023b) 0.1106 0.9941 0.1838 0.9838 0.4329 0.9055 0.6056 0.8116 0.6913 0.7571
OneForecast(ours) 0.0683 0.9976 0.1294 0.9917 0.2954 0.9577 0.4823 0.8839 0.6306 0.8025
OneForecast(Promotion) 19.11% 0.11% 19.38% 0.38% 20.66% 2.89% 13.49% 5.05% 5.59% 3.38%
Figure 3: 10-day forecast results of different models.
Figure 4: We select the latitude-weighted RMSE (lower is better) and ACC (higher is better) of several variables.
In this section, we extensively evaluate the performance of OneForecast, covering metric results, visual results, and extreme event analysis. We conduct all experiments on 128 NVIDIA A100 GPUs.
3.1 Benchmarks and Baselines
We conduct the experiments on the weatherbench2 (Rasp et al., 2024) benchmark, a subset of the fifth generation of ECMWF Reanalysis Data (ERA5) (Hersbach et al., 2020). The subset we use includes years from 1959 to 2020, which is 1959-2017 for training, 2018-2019 for validing, and 2020 for testing. We use 5 pressure level variables (each with 13 pressure levels), geopotential (Z), specific humidity (Q), temperature (T), U and V components of wind speed (U and V), and 4 surface level variables 10-meter U and V components of wind (U10M and V10M), 2-meter temperature (T2M), and Mean sea-level pressure (MSLP). For the global weather forecasts, we choose the 1.5° (121 × 240 for the global data) version of weatherbench2 as our dataset. For convenience, we just choose 120 × 240 data to train models. For regional weather forecasts, we use the original 0.25° (721 × 240 for the global data) ERA5 data. More details can be found in the Appendix B. We compare the forecast performance of OneForecast with 4 data-driven models (Pangu (Bi et al., 2023), Grapcast (Lam et al., 2023), Fengwu (Chen et al., 2023a), and Fuxi (Chen et al., 2023b)). And these models are trained using same framework and settings. We also compare extreme event prediction with IFS HRES, the state-of-the-art numerical method.
3.2 Comparison with state-of-the-art methods
We utilize two metrics, RMSE (Root Mean Square Error) and ACC (Anomalous Correlation Coefficient), to evaluate the forecast performance. More details can be found in E.1. Since the magnitudes of different variables vary greatly, we first normalize the 69 variables and then calculate the indicators for the 1300 initial conditions. As shown in Table 1, OneForecast achieves satisfactory performance compared with the state-of-the-art models. As shown in Figure 3, OneForecast are closer with the ground truth. We also show the forecast results of several important variables in Figure 4, which are not normalized. Our OneForecast performs better than other models. This improvement is primarily attributed to integrating the proposed adaptive message passing module, which enhances OneForecast’s ability to model the relationships between atmospheric states across different regions of the earth that allows for the simulation of atmospheric dynamics at various spatial and temporal scales adaptively. In summary, the forecasts exhibit more consistency with the actual physical field, effectively mitigate over-smoothing, and demonstrate superior predictive performance, particularly for extreme atmospheric values.
3.3 Regional High Resolution Forecast
Although the training cost of low-resolution forecast models is relatively low, their prediction results lack sufficient details. However, directly training on high-resolution regional data often results in poor regional forecasts performance due to issues such as missing boundary conditions and limited data samples. Although the regional forecasts method proposed by  (Oskarsson et al., 2024) improves prediction accuracy, it remains constrained by the absence of global information in high-resolution regional models, leaving room for improvement in forecast results. In contrast, the Neural Nesting Grid (NNG) method proposed in this study incorporates boundary conditions and global future information, enabling more accurate high-resolution regional predictions. Furthermore, NNG makes full use of the forecast results of global models, which achieves high-resolution regional forecasts at an exceptionally low training cost. Therefore, as shown in Figure 6, we only demonstratively conduct high-resolution predictions for two regional variables without requiring training on all variables (e.g., the 69 variables used for training global models). It can be seen that the poor long-term inference results are poor when only using regional data for training. Graph-EFM takes into account boundary conditions and the effect is improved. And our proposed NNG not only takes into account regional boundary conditions, but also makes full use of the future forecast information of the global model, which achieves stable long-term forecast performance.
Figure 5: Comparison results of 100-day forecasts between the two best models and our OneForecast.
Figure 6: High-resolution regional results. In the left figure, we select two variables, MSLP and U10M, for visualization. We compare our model with Graph-EFM and the method that directly trains on high-resolution data. The right figure shows line charts of RMSE and ACC for different models over time. These two figures demonstrate that our proposed neural nesting grid method achieves the best performance.
3.4 Extreme Events Assessment
Extreme events, such as tropical cyclones, can significantly impact human society. In this section, we evaluate our model’s ability in forecasting those extreme cyclones. As shown in Figure 1, OneForecast achieves competitive performance in typhoon tracking during two extreme events, Yagi (2018) and Molva (2020). The details of tracking algorithm can be found in Appendix E.3. Additionally, we download the forecast results of baseline models (e.g., Pangu, Fuxi and Graphcast) from weatherbench2, which is trained using high resolution (0.25°) data, to better illustrate the performance of the baselines. Although OneForecast uses lower-resolution (1.5°) data, which may limit its capacity to predict cyclones, it nevertheless shows strong forecast skills in tracking tropical cyclones comparing with the baselines.
Table 2: Comparison results between deterministic forecasts and uncertain ensemble forecast (ENS), the best results are in bold. We select 10 initial conditions, and for ENS, each results contains 50 members.
Model Metrics
4-day 10-day
RMSE ACC RMSE ACC
Pangu 0.3721 0.9334 0.6778 0.7862
Pangu (ENS) 0.3488 0.9402 0.5640 0.8236
Fengwu 0.5309 0.8913 1.1019 0.7454
Fengwu (ENS) 0.5107 0.8971 0.9288 0.7781
Graphcast 0.6897 0.7443 0.8591 0.6724
Graphcast (ENS) 0.3805 0.9378 0.6400 0.8101
Fuxi 0.4425 0.9061 0.7364 0.7556
Fuxi (ENS) 0.4053 0.9175 0.6001 0.7999
OneForecast 0.2939 0.9594 0.6153 0.8252
OneForecast (ENS) 0.3221 0.9536 0.5385 0.8446
3.5 Long-term and Ensemble Forecasts
As shown in Figure 5, we evaluate long-term forecasts with the two best models on Z500 (500hPa Geopotential). Pangu exhibits patch artifacts in 100-day forecasts, while GraphCast experiences error accumulation that degrades the forecasted physical fields, rendering them physically implausible over time. In contrast, OneForecast achieves stable long-term forecast performance, effectively capturing large-scale atmospheric states without the aforementioned issues. These results highlight OneForecast’s superior capability in maintaining accurate and physically consistent predictions over extended forecasting horizons. In Table 2, we show the results for 10 initial conditions (starting from 00:00 UTC 1 January 2020, and the interval between consecutive initial conditions is 100 time steps). The ensemble forecast (ENS) results are averaged from 50 members. Obviously, the forecasting performance is enhanced for each model when uncertainty is incorporated, and OneForecast still achieves the best performance.
3.6 Ablation Studies
To verify the effectiveness of the proposed method, as shown in Table 3, we conducted detailed ablation experiments. We introduced the following model variants: (1) OneForecast w/o Region Refined Graph, we remove the region refined mesh from the finest mesh and compute the regional metrics. (2) OneForecast w Region Refined Graph, we reserve the region refined mesh. (3) OneForecast R w/o NGG, we remove the neural nested grid (NGG) method in the regional forecasts and only use the regional data to train the model. (4) OneForecast R BF, we remove the NGG in the regional forecasts and only use the boundary forcing method to train the model. The results of the ablation experiment show that removing any module will reduce the forecast performance, further proving the rationality of the components proposed by OneForecast. (5) OneForecast R w NGG, regional forecast model with NGG. (6) OneForecast w/o MSM, we remove the multi-stream messaging module (MSM) and use a traditional MLP-based messaging module. (7) OneForecast, the full version of OneForecast for the global forecasts. For (1) and (2), we only evaluate the region-refined data. For (3), (4), and (5), we only evaluate the specific regional data. For (6) and (7), we evaluate the global data. And these results are based on 4-day forecast results. Experimental results show that whether it is a global or regional forecast task, the lack of any component will degrade the performance of OneForecast, which proves the effectiveness of the proposed method. And as shown in Figure 7, the proposed MSM can better capture of high and low frequency information, which achieves satisfactory results in long-term forecasts and extreme event forecasts, such as typhoons.
Table 3: Ablation Studies on 1.5° weatherbench2 benchmark, the best results are in bold.
Variants RMSE ACC
OneForecast w/o Region Refined Graph 0.3751 0.8954
OneForecast w Region Refined Graph 0.2736 0.9342
OneForecast R w/o NGG 0.5850 0.8405
OneForecast R BF 0.4473 0.8734
OneForecast R w NGG 0.3168 0.9325
OneForecast w/o MSM 0.3921 0.9305
OneForecast 0.2954 0.9577
Figure 7: Normalized Spectral Error of the proposed MSM and the traditional MLP-based massaging.
4 Conclusion
In this paper, we propose OneForecast, a global-regional nested weather forecasting framework leveraging multi-scale graph neural networks. By integrating dynamical systems principles with multi-grid structures, our approach refines target regions for improved capture of high-frequency features and extreme events. The adaptive information propagation mechanism, featuring dynamic gating units, mitigates over-smoothing and enhances node-edge feature representation. Additionally, the proposed neural nested grid method preserves global information for regional forecasts, effectively relieves the loss of boundary information, which improves the regional forecast performance. Empirical results show that the proposed OneForecast achieves higher prediction accuracy at global and regional scales, especially for long-term and extreme event predictions, marking a step towards more robust data-driven weather forecasting.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (42125503, 42430602).
Appendix A Proofs of Theorems
Theorem A.1 (High-pass Filtering Property of Multi-stream Messaging).
Consider an improved multi-stream message passing mechanism. Let a graph signal $\mathbf{f}\in\mathbb{R}^{N}$ have a spectrum $\hat{\mathbf{f}}=\mathbf{U}^{\top}\mathbf{f}$ under the graph Fourier basis $\mathbf{U}=[\mathbf{u}_{1},\mathbf{u}_{2},\ldots,\mathbf{u}_{N}]$, where $\mathbf{L}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$ is the normalized graph Laplacian and $\boldsymbol{\Lambda}=\mathrm{diag}(\lambda_{1},\lambda_{2},\ldots,\lambda_{N})$ with $0\leq\lambda_{1}\leq\cdots\leq\lambda_{N}\leq 2$ Define the frequency response function of the message passing operator by $\rho:\lambda\mapsto\mathbb{R}.$ If the dynamic gating weights satisfy
$g^{(h,e)}_{i},\;g^{(h,s)}_{i},\;g^{(h,d)}_{i}\;\propto\;\bigl{|}\lambda_{i}-1\bigr{|}\;+\;\epsilon\quad(\epsilon>0),$ (20)
then there exist constants $\alpha>0$ and $\kappa>0$ such that the frequency response $\rho(\lambda_{i})$ of the operator satisfies
$\rho(\lambda_{i})\;\geq\;\alpha\,\bigl{|}\lambda_{i}-1\bigr{|}\quad\text{and}\quad\rho(\lambda_{i})\;\geq\;\kappa\,\lambda_{i},$ (21)
which means the operator is a strictly high-pass filter.
Proof.
We consider a multi-stream message passing operator that we denote by $\mathbf{G}$. This operator depends on both node features and dynamic gating on edges. We let $\mathbf{f}\in\mathbb{R}^{N}$ be an arbitrary graph signal, and we write its graph Fourier transform as $\hat{\mathbf{f}}=\mathbf{U}^{\top}\mathbf{f}$, where $\mathbf{L}=\mathbf{U}\boldsymbol{\Lambda}\mathbf{U}^{\top}$ is the normalized Laplacian and $\boldsymbol{\Lambda}=\mathrm{diag}(\lambda_{1},\ldots,\lambda_{N})$ with $0\leq\lambda_{1}\leq\cdots\leq\lambda_{N}\leq 2$.
Step 1: Integral analogy of gating on a discrete graph. We first recall that on a continuous domain, a gating operator often admits a representation of the form
[…]
where $\mathcal{K}(x,\xi)$ is a kernel that encodes the gating weights. On a discrete graph, the integral turns into a finite sum. Hence for node $i$, we write
$(\mathbf{G}\mathbf{f})_{i}\;=\;\sum_{j\in\mathcal{N}_{i}}\mathcal{K}_{ij}\,f_{j},$
where $\mathcal{N}_{i}$ denotes the neighbors of node $i$, and $\mathcal{K}_{ij}$ depends on the dynamic gating parameters $\bigl{(}g^{(h,e)}_{i},g^{(h,s)}_{i},g^{(h,d)}_{i}\bigr{)}$. We assume these gating parameters scale proportionally to $\bigl{|}\lambda-1\bigr{|}+\epsilon$, which implies larger weights when $\lambda$ is around high or mid-frequency regions.
Step 2: Spectral decomposition of the operator. We decompose $\mathbf{f}$ in the eigenbasis of $\mathbf{L}$:
$\mathbf{f}\;=\;\sum_{\ell=1}^{N}\hat{f}_{\ell}\,\mathbf{u}_{\ell},\quad f_{j}\;=\;\sum_{\ell=1}^{N}\hat{f}_{\ell}\,(\mathbf{u}_{\ell})_{j}.$
The operator $\mathbf{G}$ acts on $\mathbf{u}_{\ell}$ with some gain factor $\rho(\lambda_{\ell})$, which we call the frequency response. In other words, we write
$\mathbf{G}\,\mathbf{u}_{\ell}\;=\;\rho(\lambda_{\ell})\,\mathbf{u}_{\ell}.$
Thus the value of $\rho(\lambda_{\ell})$ reveals how the operator scales the amplitude of the $\ell$-th eigenmode.
Step 3: High-pass filtering behavior from gating design. We now analyze the effect of gating weights $g^{(h,e)}_{i},\,g^{(h,s)}_{i},\,g^{(h,d)}_{i}$ that satisfy
$g^{(h,e)}_{i},\,g^{(h,s)}_{i},\,g^{(h,d)}_{i}\;\propto\;|\lambda_{i}-1|+\epsilon.$
Because $\lambda_{i}\in[0,2]$, when $\lambda_{i}$ is close to $2$, it represents a high-frequency component on the graph. In that regime, the gating weights become larger, and the message passing operator $\mathbf{G}$ amplifies those components. Similarly, when $\lambda_{i}$ is near $1$, the factor $|\lambda_{i}-1|$ can still be significant enough to enhance mid-to-high frequencies. Conversely, for $\lambda_{i}$ near 00 (low frequency), the gating is relatively small and thus tends to suppress those components.
Step 4: Combining inequalities to show strictly high-pass. We combine partial inequalities for different ranges of $\lambda_{i}$. Since $0\leq\lambda_{i}\leq 2$, we use the gating assumption to show there are positive constants $\alpha$ and $\kappa$ such that
$\rho(\lambda_{i})\;\geq\;\alpha\,\bigl{|}\lambda_{i}-1\bigr{|}\quad\text{and}\quad\rho(\lambda_{i})\;\geq\;\kappa\,\lambda_{i}.$
Hence the operator $\mathbf{G}$ behaves like a high-pass filter, because it provides higher gain to higher (or mid-high) frequency components and less gain to low-frequency components. Therefore, $\mathbf{G}$ is a strictly high-pass operator in the graph spectral domain.
This completes the proof of Theorem A.1. ∎
Appendix B Data Details
B.1 Dataset
In this section, we introduce the dataset used in this paper detailedly. For the global forecasting, we conduct experiments on the weatherbench2 (Rasp et al., 2024) benchmark, a subset of ERA5 reanalysis data (Hersbach et al., 2020). The weatherbench2 benchmark we used is the version of 1.5° resolution (121 × 240), which spans from 1959 to 2020. This subset contains 5 variables (Z, Q, T, U, V) with 13 pressure levels (50 hPa, 100 hPa, 150 hPa, 200 hPa, 250 hPa, 300 hPa, 400 hPa, 500 hPa, 600 hPa, 700 hPa, 850 hPa, 925 hPa and 1,000 hPa) and 4 variables (U10M, V10M, T2M, MSLP) with surface level. For the regional forecasting, a higher resolution data (0.25° resolution) of ERA5 is also used, which can be downloaded from https://cds.climate.copernicus.eu/, the official website of Climate Data Store (CDS). All the data we used are shown in Table 4. For both global and regional forecasts, we use the data from 1959 to 2017 for training, 2018 to 2019 for validating, and 2020 for testing.
Table 4: The data details in this work.
Task
Variable
Name
Layers
Spatial
Resolution
Dt
Lat-Lon
Range
Time
Global Geopotential (Z) 13 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
Specific Humidity (Q) 13 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
Temperature (T) 13 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
U Component of Wind (U) 13 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
V Component of Wind (V) 13 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
10 Metre U Wind Component (U10M) 1 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
10 Metre V Wind Component (V10M) 1 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
2 Metre Temperature (T2M) 1 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
Mean Sea Level Pressure (MSLP) 1 1.5° 6h -90°S-180°Wsimilar-$\sim$90°N180°E 1959$\sim$2020
Regional Mean Sea Level Pressure (MSLP) 1 0.25° 6h 7.5°W114°Esimilar-$\sim$ 36°W172.5°E 1959$\sim$2020
10 Metre U Wind Component (U10M) 1 0.25° 6h 7.5°W114°Esimilar-$\sim$ 36°W172.5°E 1959$\sim$2020
B.2 Data preprocessing
Different atmospheric variables have large differences in numerical magnitude. To allow the model focusing on predictions rather than learning the differences between variables, we normalized the data before putting the data into the model. We calculated the mean and standard deviation of all variables using data from 1959 to 2017 (training set). Each variable has a corresponding mean and standard deviation. Before putting the data into the model, we subtracted the change from the respective mean and divided it by the standard deviation.
Appendix C Algorithm
We summarize the overall framework of OneForecast in Algorithm 1.
Algorithm 1 OneForecast Framework for Global Weather Forecast
0:  Initial atmospheric condition $Z_{t}$.
0:  Next step atmospheric state $Z_{t+1}$.
1:  Initialize OneForecast
2:  repeat
3:     Encoder
4:     Embedding features of grid nodes $Z_{t}$, mesh nodes $\mathcal{V}$, mesh edges $\mathcal{E}$, grid to mesh edges $\mathcal{E}^{\mathrm{G}2\mathrm{M}}$, and mesh to grid edges $\mathcal{E}^{\mathrm{M}2\mathrm{G}}$ into latent space using respective MLP: ($\mathcal{V}_{f}^{G}$, $h$, $\mathcal{E}_{f}$, $\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}$, $\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f}$) = MLPs($Z_{t}$, $\mathcal{V}$, $\mathcal{E}$, $\mathcal{E}^{\mathrm{G}2\mathrm{M}}$, $\mathcal{E}^{\mathrm{M}2\mathrm{G}}$)
5:     Project the atmospheric state from the lat-lon grid into the mesh nodes: ${\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime}=\mathrm{ESMLP}(\mathcal{V}_{f}^{G},h,\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f})$, $h^{\prime}=\mathrm{MLP}_{e1}(h,\sum{\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime})$
6:     Update grid node feature: ${\mathcal{V}_{f}^{G}}^{\prime}=\mathrm{MLP}_{e2}(\mathcal{V}_{f}^{G})$
7:     Apply residual connection to update the feature of grid to mesh edge, mesh node, and grid node again: ${\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}={\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime}+{\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}$, $h^{\prime}=h^{\prime}+h$, ${\mathcal{V}_{f}^{G}}={\mathcal{V}_{f}^{G}}^{\prime}+{\mathcal{V}_{f}^{G}}$
8:     Multi-stream Messaging
9:     Apply dynamic multi-head gated edge update module (DMG) to update edge feature: $\mathcal{E}_{f}^{\prime}=DMG(\mathcal{E}_{f},h_{s},h_{r})$
10:     Apply multi-head node attention mechanism (MHA) to update mesh node feature: $h^{\prime}=MHA(h,\sum{\mathcal{E}_{f}}^{\prime})$
11:     Apply residual connection to update the feature of edge and mesh node: $\mathcal{E}_{f}=\mathcal{E}_{f}^{\prime}+\mathcal{E}_{f}$, $h=h^{\prime}+h$
12:     Decoder Project the feature from mesh back to lat-lon grid: ${\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f}}^{\prime}=\mathrm{ESMLP}(\mathcal{V}_{f}^{G},\mathcal{E}_{f},\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f})$, ${\mathcal{V}_{f}^{G}}^{\prime}={\mathrm{MLP}_{d1}}(\mathcal{V}_{f}^{G},\sum{\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f}}^{\prime})$, $\mathcal{V}_{f}^{G}={\mathrm{MLP}}_{d2}(\mathcal{V}_{f}^{G^{\prime}}+\mathcal{V}_{f}^{G})$, $Z_{t+1}=Z_{t}+{\mathcal{V}_{f}^{G}}$
13:  until converged
14:  return $OneForecast$
Appendix D Model Details for Global Forecasts
D.1 Earth-specific Region Refined Multi-scale Graph
The graph used in OneForecast can be defined as: $\mathcal{G}(\mathcal{V}^{G},\mathcal{V},\mathcal{E},\mathcal{E}^{\mathrm{G}2\mathrm{M}}$, $\mathcal{E}^{\mathrm{M}2\mathrm{G}})$.
Grid Nodes. $\mathcal{V}^{G}$ is the ensemble of grid nodes, which contains $120\times 240=28800$ nodes for 1.5° global data in global forecast task. And each node consists of 69 atmospheric features (5 variables at 13 pressure levels and 4 variables at surface level, $5\times 13+4=69$). Since we just consider 1 step historical state, the input features of OneForecast are 69. For regional forecast, the region size can be arbitrary within the permission of GPU memory. For simplicity, we choose the region size of $120\times 240$ from 0.25° data, the node is still $120\times 240=28800$.
Mesh Nodes. $\mathcal{V}$ is the ensemble of mesh nodes, which contains multi-scale mesh nodes of different fineness and region refined mesh nodes that cover the global area. The mesh nodes are distributed over a refined icosahedron that has undergone five levels of subdivision, and the coarsest icosahedron consists of 12 vertices and 20 triangular faces. By dividing each triangular face into four smaller triangles, an additional node is generated at the midpoint of each edge. The new nodes are then projected back onto the unit sphere, gradually refining the grid. To enhance the forecasting performance in key regions, we further refine specific areas of the finest mesh, achieving localized mesh densification. For the global forecast task, we refine the 2 areas: 0°N105°Esimilar-$\sim$30°N160°E and 10°N-95°Wsimilar-$\sim$30°N-35°W. The features of each node include the cosine value of the latitude, as well as the sine and cosine values of the longitude. We only keep the finest mesh nodes, since the nodes on the coarse mesh are its subset. In total, the graph structure of OneForecast comprises 12337 mesh nodes, each characterized by three features.
Mesh Edges. $\mathcal{E}$ are the bidirectional edges that connect mesh nodes (sender and receiver nodes). Similar to mesh nodes, there are corresponding edges for each scale of mesh, and $\mathcal{E}$ is the ensemble of multi-scale edges. And the features of each edge include the length of edge, the 3D position difference between sender and receiver nodes. In total, OneForecast comprises 98296 mesh edges, each characterized by four features.
Grid2Mesh Edges. $\mathcal{E}^{\mathrm{G}2\mathrm{M}}$ are the the unidirectional edges that used in the encoder, which connect grid and mesh nodes. To ensure that each grid node has a corresponding mesh node connected to it, we add $\mathcal{E}^{\mathrm{G}2\mathrm{M}}$ to grid nodes and mesh nodes if the distance between them is less than or equal to 0.6 times the edge length of the finest $\mathcal{E}$. Similar to mesh edge $\mathcal{E}$, each grid2mesh edge comprises 4 features, and OneForecast has 49233 grid2mesh edges in total.
Mesh2Grid Edges. $\mathcal{E}^{\mathrm{M}2\mathrm{G}}$ are the unidirectional edges that used in the decoder, which connect grid and mesh nodes. For each grid node, we find the triangle face containing it on the finest mesh and connect 3 mesh nodes to it. Similar to other edges, each mesh2grid edge has 4 features. In total, OneForecast has 86,400 mesh2grid edges.
D.2 Encoder
We first apply embedder MLP to map the data to the latent space, which can be defined as:
$MLP=LN(SiLU(Linear(x_{embedder}))),$ (22)
$(\mathcal{V}_{f}^{G},h,\mathcal{E}_{f},\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f},\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f})=MLPs(Z_{t},\mathcal{V},\mathcal{E},\mathcal{E}^{\mathrm{G}2\mathrm{M}},\mathcal{E}^{\mathrm{M}2\mathrm{G}}),$ (23)
where, $x_{embedder}$ is the input of embedder MLP. For the linear function, we set the latent dim to 512. $SiLU(\cdot)$ is the SiLU activation function, $LN(\cdot)$ is the layernorm function. $Z_{t}$, $\mathcal{V}$, $\mathcal{E}$, $\mathcal{E}^{\mathrm{G}2\mathrm{M}}$, and $\mathcal{E}^{\mathrm{M}2\mathrm{G}}$ are embedded features of grid nodes, mesh nodes, mesh edges, grid to mesh edges, and mesh to grid edges.
We then project the atmospheric state from the lat-lon grid into the mesh nodes. Specifically, we first update the edge features through an Edge Sum MLP (ESMLP):
$\mathcal{E}_{f}^{\mathrm{G}2\mathrm{M}\prime}=\mathbf{W}_{e}\mathcal{E}_{f}^{\mathrm{G}2\mathrm{M}}$ (24)
$\mathcal{V}_{f}^{G\prime}=\mathbf{W}_{s}\mathcal{V}_{f}^{G},$ (25)
$h^{\prime}=\mathbf{W}_{d}h+\mathbf{b}_{d},$ (26)
$\mathbf{h}_{\mathrm{sum}}=\mathcal{E}_{f}^{\mathrm{G}2\mathrm{M}\prime}+\mathcal{V}_{f}^{G\prime}+h^{\prime},$ (27)
${\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime}=\operatorname{LN}\left(\mathbf{W}_{\mathrm{ESMLP}}\sigma\left(h_{out}\right)+\mathbf{b}_{\mathrm{ESMLP}}\right),$ (28)
where, $\mathbf{W}_{e}$, $\mathbf{W}_{s}$, $\mathbf{W}_{d}$ are the linear transformation matrix of grid2mesh edge features, grid node feature, and mesh node features. $\mathbf{W}_{\text{ESMLP }}$ is the linear transformation matrix of output layer. $b_{d}$ is the bias of mesh node during linear transformation, $b_{ESMLP}$ is the bias vector of ESMLP. In summary, the grid2mesh edge update process can be define as:
${\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime}=\mathrm{ESMLP}(\mathcal{V}_{f}^{G},h,\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}).$ (29)
After updating the grid2node features, we update the mesh node features using another MLP:
$h^{\prime}=\mathrm{MLP_{e1}}(h,\sum{\mathcal{E}^{\mathrm{G}2\mathrm{M}}_{f}}^{\prime}),$ (30)
where, $\sum\mathcal{E}_{f}^{\mathrm{G}2\mathrm{M}^{\prime}}$ are the edges that arrives at mesh node.
Then, we update the grid node features using another MLP:
$\mathcal{V}_{f}^{G^{\prime}}=\operatorname{MLP}_{e2}\left(\mathcal{V}_{f}^{G}\right).$ (31)
Finally, residual connections are applied to update the feature of grid to mesh edge, mesh node, and grid node again.
D.3 Multi-stream Messaging
The proposed multi-stream messaging is implemented by an adaptive messaging mechanism, which contains a dynamic multi-head gated edge update module and a multi-head node attention module. This part has been introduced in detail in the main text, so we only added the hyperparameter settings here. For the dynamic multi-head gated edge update module, the dimensions of the gating vector are set to 64. In the multi-head node attention module, the $MLP_{a}$ used to calculate the attention score consists of a linear layer, a SiLU activation function, a linear layer, and a Sigmoid function. The hidden dimension of the linear layer is 64.
D.4 Decoder
In the decoder, we map the feature from mesh back to lat-lon grids, similar to encoder, we first up the mesh2grid features:
$\mathcal{E}_{f}^{\mathrm{M}2\mathrm{G}^{\prime}}=\operatorname{ESMLP}\left(\mathcal{V}_{f}^{G},\mathcal{E}_{f},\mathcal{E}_{f}^{\mathrm{M}2\mathrm{G}}\right).$ (32)
Then, we update the grid node features:
${\mathcal{V}_{f}^{G}}^{\prime}={\mathrm{MLP}}_{d1}(\mathcal{V}_{f}^{G},\sum{\mathcal{E}^{\mathrm{M}2\mathrm{G}}_{f}}^{\prime}),$ (33)
where, $\sum\mathcal{E}_{f}^{\mathrm{M}2\mathrm{G}^{\prime}}$ are the edges that arrives at grid node.
After that, a residual connection is applied to update the grid node features again:
$\mathcal{V}_{f}^{G^{\prime}}=\mathcal{V}_{f}^{G^{\prime}}+\mathcal{V}_{f}^{G}.$ (34)
Finally, we apply a MLP to predict the next step results:
$\mathcal{V}_{f}^{G}=\operatorname{MLP}_{d2}\left(\mathcal{V}_{f}^{G^{\prime}}+\mathcal{V}_{f}^{G}\right).$ (35)
Appendix E Experiments Details
E.1 Evaluation Metric
We utilize two metrics, RMSE (Root Mean Square Error) and ACC (Anomalous Correlation Coefficient), to evaluate the forecasting performance, which can be defined as:
$\small{RMSE}(k,t)=\sqrt{\frac{\sum\limits_{i=1}^{N_{\text{lat}}}\sum\limits_{j=1}^{N_{\text{lon}}}L(i)\left(\hat{\mathbf{A}}_{ij,t}^{k}-\mathbf{A}_{ij,t}^{k}\right)^{2}}{N_{\text{lat}}\times N_{\text{lon}}}},$ (36)
$\small\operatorname{ACC}(k,t)=\sqrt{\frac{\sum\limits_{i=1}^{N_{\text{lat}}}\sum\limits_{j=1}^{N_{\text{lon}}}L(i)\hat{\mathbf{A^{\prime}}}_{ij,t}^{k}\mathbf{A^{\prime}}_{ij,t}^{k}}{\sum\limits_{i=1}^{N_{\text{lat}}}\sum\limits_{j=1}^{N_{\text{lon}}}L(i)\left(\hat{\mathbf{A^{\prime}}}_{ij,t}^{k}\right)^{2}\times\sum\limits_{i=1}^{N_{\text{lat}}}\sum\limits_{j=1}^{N_{\text{lon}}}L(i)\left(\mathbf{A^{\prime}}_{ij,t}^{k}\right)^{2}}},$ (37)
where $\mathbf{A}_{i,j,t}^{v}$ represents the value of variable v at horizontal coordinate $(i,j)$ and time t. Latitude-dependent weights are defined as $L(i)=N_{\text{lat }}\times\frac{\cos\phi_{i}}{\sum_{i’=1}^{N_{\text{lat}}}\cos\phi_{i’}}$, where $\phi_{i}$ is the latitude at index i. The anomaly of $A$, denoted as $A^{\prime}$, is computed as the deviation from its climatology, which corresponds to the long-term mean of the meteorological state estimated from 59 years of training data. To evaluate model performance, RMSE and ACC are averaged across all time steps and spatial coordinates, providing summary statistics for variable $v$ a given lead time $\Delta t$.
E.2 Model Training
We train all baseline models and OneForecast using the same training framework. We set the total model training epochs to 110, the initial learning rate is 1e-3, and use the cosine annealing scheduler to adjust the learning rate until the model converged. The model codes of Pangu, Graphcast, and Fengwu we used are released by NVIDIA modulus (https://github.com/NVIDIA/modulus). The model code of Fuxi is obtained by sending an email to the author. For all models, we select the checkpoint that performed best on the validation set for comparative analysis.
E.3 Typhoon Tracking
To track the eye of a tropical cyclone, we follow (Bi et al., 2023; Magnusson et al., 2021) to find the local minimum of mean sea level pressure (MSLP). The time step of forecast lead time is set to be 6 hours. Specifically, once the initial position of the cyclone eye is provided, we iteratively search for a local minimum of MSLP that meets the following criteria:
- There is a maximum 850hPa relative vorticity greater than $5\times 10^{-5}$ within a 278km radius (in the Northern Hemisphere).
- There is a maximum thickness between 850hPa and 200hPa within a 278km radius when the cyclone is extratropical.
- The maximum 10m wind speed exceeds 8m/s within a 278km radius when the cyclone is over land.
Once the cyclone eye is identified, the tracking algorithm continues to find the next position within a 445km vicinity.
This study focuses on two extreme cyclones: Tropical Storm Yagi and Severe Typhoon Molave. The Yagi formed near Ito To, Japan on August 6, 2018, and landed over Wenling, China on August 12. The Molave formed on October 11, 2020, and landed over the Philippines on October 25, 2020. The initial conditions for these two cyclones are set at 0:00 UTC, August 6, 2018 and 0:00 UTC, 11 October 2020, respectively. Since there is no Fuxi results in 2018 on weatherbench2, we can not compare it with Yagi.
Appendix F Related Work
F.1 Deep Learning based Weather Forecasting
Global Weather Forecasting. Global weather forecasting has seen significant progress with deep learning models. FourCastNet, based on Fourier neural operators, provides global forecasts comparable to traditional numerical methods like IFS, but at much higher speeds (Pathak et al., 2022). Pangu, utilizing the Swin Transformer, exceeds NWP methods, incorporating earth-specific location embeddings for better performance (Bi et al., 2023). The Spherical Fourier Neural Operator (SFNO) extends Fourier methods using spherical harmonics, offering more stable long-term predictions (Bonev et al., 2023). FuXi focuses on long-term forecasting, achieving a 15-day forecasts comparable to ECMWF (Chen et al., 2023b). GraphCast leverages message-passing networks to improve efficiency and forecasting accuracy (Lam et al., 2023), and GenCast builds on this to enhance ensemble forecasting (Price et al., 2023). Further, diffusion models like those in (Li et al., 2024) generate probabilistic ensembles by sampling, while NeuralGCM (Kochkov et al., 2024) focuses on atmospheric circulation with a dynamic core, offering climate simulation capabilities but at higher training and inference costs.
Regional Weather Forecasting. The goal of regional weather forecasting is to enhance local prediction accuracy with high-resolution models. CorrDiff (Mardani et al., 2023) combines U-Net and diffusion models to improve local forecasts. MetaWeather (Kim et al., 2024) adapts global forecasts to regional contexts using meta-learning. GNNs are also widely applied in regional forecasting, with Graphcast (Lam et al., 2023) enhancing accuracy by modeling complex spatial dependencies. MetNet-3 (Espeholt et al., 2022) offers high-accuracy forecasts for weather variables, such as precipitation, temperature, and wind speed, at 2-minute intervals and 1–4 km resolution, outperforming traditional models like HRRR. NowcastNet (Zhang et al., 2023) and DGMR (Ravuri et al., 2021) excel in short-term extreme precipitation forecasts using deep generative models and radar data. In spatiotemporal prediction, NMO (Wu et al., 2024b) models the evolution of physical dynamics, providing new insights for local weather forecasting. Similarly, SimVP (Gao et al., 2022) and PastNet (Wu et al., 2024c) achieve good results in forecasting local precipitation evolution using spatiotemporal convolution methods.
F.2 Numerical analysis methods
Multigrid methods (McCormick, 1987; Wesseling, 1995; Hackbusch, 2013; Bramble, 2019; Hiptmair, 1998; Brandt et al., 1983; Borzi & Schulz, 2009) and nested grid strategies (Miyakoda & Rosati, 1977; Zhang et al., 2012; Sullivan et al., 1996) are widely used to solve PDEs and handle multi-scale problems (Debreu & Blayo, 2008; Xue et al., 2000). Multigrid methods use grids of different resolutions to transfer information and accelerate iterations. They efficiently solve large-scale problems and improve computational accuracy. By eliminating low-frequency errors on coarse grids and high-frequency errors on fine grids, multigrid methods effectively handle error convergence at different scales (He & Xu, 2019; He et al., 2023; Shao et al., 2022). Nested grid strategies embed higher-resolution fine grids into regions of interest based on a global coarse grid to capture local complex physical phenomena in detail. In weather forecasting, this method provides large-scale background fields on a global scale while refining the grid for target regions to accurately simulate the evolution of local weather systems and the occurrence of extreme events (Bacon et al., 2000).
Appendix G Additional Results
We present more additional results in Figure 8, 9, 10 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, including 18 variables that are importmant to weather forecasting, each with results ranging from 6 hours to 10 days. These additional results further demonstrate the effectiveness of OneForecast. Same as the Figure 3 , the initial conditions is 00:00 UTC, 1 January 2020.
Figure 8: 6-hour forecast results of different models.
Figure 9: 0.5-day forecast results of different models.
Figure 10: 1-day forecast results of different models.
Figure 11: 1.5-day forecast results of different models.
Figure 12: 2-day forecast results of different models.
Figure 13: 2.5-day forecast results of different models.
Figure 14: 3-day forecast results of different models.
Figure 15: 3.5-day forecast results of different models.
Figure 16: 4-day forecast results of different models.
Figure 17: 4.5-day forecast results of different models.
Figure 18: 5.0-day forecast results of different models.
Figure 19: 5.5-day forecast results of different models.
Figure 20: 6.0-day forecast results of different models.
Figure 21: 6.5-day forecast results of different models.
Figure 22: 7.0-day forecast results of different models.
Figure 23: 7.5-day forecast results of different models.
Figure 24: 8.0-day forecast results of different models.
Figure 25: 8.5-day forecast results of different models.
Figure 26: 9.0-day forecast results of different models.
Figure 27: 9.5-day forecast results of different models.
Figure 28: 10.0-day forecast results of different models.